### v1

# Revisiting Gamma-Ray Emission of the Supernova Remnant RCW 103

Abstract
We analyze more than 15 years of the -ray data, obtained with the Large Area Telescope (LAT) onboard the Fermi Gamma-ray Space Telescope (Fermi), for the region of the young supernova remnant (SNR) RCW 103, since the nearby source 4FGL J1616.25054e, counterpart to HESS J1616518 and 13 arcmin away from the SNR, is determined to be extended in the more recent Fermi-LAT source catalog. Different templates for 4FGL J1616.25054e and RCW 103 are tested, and we find that a point source with a power-law (PL) spectrum at the southern limb of the SNR best describes the detected -ray emission. The photon index of the PL emission is , softer than the previously reported when the counterpart to HESS J1616518 was considered to be a point source (which likely caused mis-identification of extended emission at RCW 103). In order to produce the -ray emission in a hadronic scenario, we estimate that protons with an index2.4 PL energy distribution are needed. These results fit with those from multi-wavelength observations that have indicated the remnant at the southern limb is interacting with a molecular cloud.
1 Introductions
The supernova remnant (SNR) RCW 103 (or G332.40.4; Rodgers et al. 1960) has been extensively studied at multi-wavelengths. Its role as a valuable sample of a young SNR, as well as its product, 1E 1613485055 as a likely magnetar (or previously known as a central compact object; Tuohy & Garmire 1980; D’Aì et al. 2016; Rea et al. 2016), have attracted much attention to it. Adopting its most recently derived distance of 3.1 kpc (Reynoso et al., 2004), its age has been estimated to be 2 kyr (Carter et al., 1997) or 4 kyr (Braun et al., 2019; Leahy et al., 2020). The SNR appears in the sky as a nearly circular, shell-like structure, with a diameter of 10′. Notably, its southern part is brighter than the regions of other directions (Figure 1), which has been seen at multi-wavelengths from radio frequencies to X-rays (e.g., Pinheiro Gonçalves et al. 2011).
SNRs are primary -ray sources in the Galaxy (e.g., Zeng et al. 2019; Ballet et al. 2023). Cosmic-ray particles are produced from their shock fronts through the diffuse shock acceleration process (e.g., Drury 1983). The particles emit -rays via the hadronic or/and leptonic processes; the former involves proton-proton collisions and the latter electron radiation processes including synchrotron and inverse Compton scattering (ICS).
Previously, Xing et al. (2014) has reported likely detection of extended -ray emission from RCW 103 in the 1–300 GeV data obtained with the Large Area Telescope (LAT) on board the Fermi Gamma-ray Space Telescope (Fermi). At the time, only 5 years of the Fermi-LAT data were collected and the source model for the data analysis was based on the Fermi-LAT second source catalog (Nolan et al., 2012). In addition, the SNR is located in a crowded region, with two very-high-energy (VHE) sources nearby. They are HESS J1616508 and HESS J1614518, which were detected in the High Energy Stereoscopic System (HESS) survey (Aharonian et al., 2005, 2006; H. E. S. S. Collaboration et al., 2018). In particular, the first source is 13 arcmin away from RCW 103, considered to be very close given the spatial resolutions of Fermi-LAT and VHE detectors. In the previous analysis, the GeV counterpart to HESS J1616508 was taken as a point source (2FGL J1615.05051), but according to the latest Fermi LAT source catalog (4FGL-DR4; Ballet et al. 2023), both GeV counterparts to HESS J1616508 and HESS J1614518 are extended sources, named as 4FGL J1616.25054e and 4FGL J1615.35146e respectively (see Figure 1). With the updated source information taken into consideration, we re-analyzed 15 years of the Fermi-LAT data for RCW 103, and obtained new results. In this paper, we report the results.
2 Data Analysis and Results
2.1 FERMI-LAT data and source model
We selected the LAT events in energy range of from 0.1 to 500 GeV from the updated Fermi LAT Pass 8 database (P8R3). The region of interest (RoI) was set to be centered at HESS J1616508 with a size of 20∘ 20∘. The start and end time of the events were August 4 2008, 15:43:39 (UTC) and December 11 2023, 23:26:35 (UTC), respectively. We included the events in the SOURCE event class with zenith angles less than 90 degree to avoid the Earth’s limb contamination, and excluded events with ‘bad’ quality flags. Both selections are recommended by the LAT team111http://fermi.gsfc.nasa.gov/ssc/data/analysis/scitools/.
A source model was constructed by including all sources within a 20-degree radius of HESS J1616508 based on 4FGL-DR4 (Ballet et al., 2023). The positions and spectral parameters for these sources provided in the catalog were adopted. For sources within 5-degree radius of HESS J1616508, we set their spectral parameters free, and for the remaining sources, their spectral parameters were fixed at their catalog values. In this latest source catalog, the counterpart of HESS J1616508 is an extended source described by a uniform disk with a radius of 032, and its -ray spectrum is modeled with a power law (PL). The spectral model gll_iem_v07.fit and spectral file iso_P8R3_SOURCE_V3_v1.txt were included as the Galactic and extragalactic diffuse emission, respectively. The normalizations of the two diffuse-emission components were always set as free parameters in our analysis.
In constructing 4FGL catalogs (Abdollahi et al., 2020), the LAT events below 316 MeV were given low weights, especially in the vicinity of the Galactic plane. In addition, only Front events were utilized in the low energy band below 316 MeV (Abdollahi et al., 2020) in order to avoid the relatively poor point spread function (PSF) and the contamination from the Earth limb. Given these, we did not include the events below 300 MeV in our analysis since the target is located in the Galactic plane with a few nearby sources around.
2.2 Maximum Likelihood Analysis
With the source model, we performed the standard binned likelihood analysis to the LAT data in 0.3–500 GeV. A PL photon index =2.020.03 and a 0.3–500 GeV photon flux photon cm-2 s-1 were obtained for HESS J1616508, and the Test Statistic (TS) value was 1640.
To examine the source field, we calculated TS maps centered at HESS J1616508 in different energy ranges. All sources in the source model, except
the counterpart of HESS J1616508, were included and thus removed in the TS maps.
In Figure 1, we showed TS maps in the energy ranges of
0.3–500 GeV and 10–500 GeV as the representative ones, where the latter
helps reveal the emission in a high, spatial resolution222http://www.slac.stanford.edu/exp/glast/groups/canda
/lat_Perfomance.htm
(smaller than the separation between HESS J1616508 and RCW 103).
As can be seen in the low-energy TS map, possibly extended emission within
the source region of HESS J1616508 (H. E. S. S. Collaboration et al., 2018) is detected, which is
enclosed by the extended source 4FGL J1616.25054e given in 4FGL-DR4.
We also noted that the other
extended source 4FGL J1615.35146e, corresponding to HESS J1614518, is
cleanly removed. However, the position of the highest TS values within
4FGL J1616.25054e is off the center towards the south-east direction, where
RCW 103 is located.
In the high-energy TS map, the high TS-value region is possibly composed of
two components, one slightly off the center of 4FGL J1616.25054e and
the other close to the location of RCW 103. Thus, it is likely that emission
from RCW 103 was detected even though it is close to an extended source.
| Source Model | 2) |
| HESS0.32 | 0 |
| HESS0.40 | 43.2 |
| CCO HESS0.32 | 57.1 |
| Radiodisk HESS0.32 | 63.5 |
| Radioprofile HESS0.32 | 71.3 |
| Radiops HESS0.32 | 82.1 |
| Radiops HESS0.38 | 106.4 |
2.3 Spatial Distribution Analysis and Source Identification
In order to fully probe the -ray emission in the region, we ran a series of spatial analysis with different spatial templates considered. The base model was that given in 4FGL-DR4, which sets a 032 uniform disk template for the GeV emission of HESS J1616508 (HESS0.32 model). All the likelihood results were compared to the likelihood of this base model.
We first tested a uniform disk with radius ranging from 002 to 050 (with a step of 002) centered at HESS J1616508 as a check for the source’s GeV -ray emission. A simple PL was used to model the emission. We found that when radius was 040 (HESS0.40 model), a maximum value was obtained, which indicates a fit improvement at a 6.6 significance level (calculated from ; Table 1) compared to the HESS0.32 model.
We then added different model templates in the source model at the position
of the central compact object (CCO) of RCW 103. The templates included
a point source (PS; CCO model), a uniform disk with a radius of
5 (given RCW 103’s size of 10 in diameter detected at
radio frequencies; Radiodisk model), and an emission template
derived from
the Molonglo Observatory Synthesis Telescope (MOST) radio map333http://snrcat.physics.umanitoba.ca/
SNRrecord.php?id=G332.4m00.4
(Radioprofile model). For HESS J1616508, the catalog
HESS0.32 template was still used. Performing the maximum likelihood
analysis to the data with each of the models, CCO HESS0.32,
Radiodisk HESS0.32, and
Radioprofile HESS0.32,
the likelihood values were obtained and compared to that of the base model
(see Table 1). The model of Radioprofile HESS0.32
provided the best fit,
as it induced a 8.4 fit improvement compared to only HESS0.32.
In addition, we noted that the fit improvement was 5.3 when
the non-catalog model HESS0.40 we obtained above was considered for
comparison.
These results indicate that extra emission is present at the position
of RCW 103.
We further removed the counterpart of HESS J1616508 from TS maps to reveal the residual emission in the RCW 103 region. Two models for the HESS J1616508 source were considered, one the HESS0.32 model and the other a uniform disk with a radius of 038 (HESS0.38 model). The latter was used because of the results obtained in the following analysis. As a check, we first calculated two TS maps in 0.3–5 GeV respectively with the two models and verified that no significant differences between the two models were found. We then calculated TS maps in 5–500 GeV without and with HESS0.38 being removed (Figure 2). Comparing the two TS maps, it is clear to see the removal of the emission from HESS J1616508 and the presence of the emission at RCW 103. However, the latter did not exactly match the SNR’s circular region, suggesting further analysis investigation would be needed.
We ran gtfindsrc in the Fermitools to the 5 GeV data and determined the position of the residual emission. We obtained R.A.=244427, Decl.=51119 (equinox J2000.0) with a 1 nominal uncertainty of 0047. The CCO is 0081 away from this position and outside of the 2 error circle (0077). Matching the position and 1/2 error circles to the MOST radio map of RCW 103 (Figure 1), it can be seen that the source region is actually over the brightest part of the SNR’s radio shell. Given this, we set a model for the residual emission which is at the position of the brightest part of the shell (Radiops model), and performed the likelihood analysis. The used -ray template for HESS J1616508 was HESS0.32. Comparing the values, we found that the residual emission is more likely a point source located at the bright radio shell, as the fit improvement of the model Radiops HESS0.32 is at 4.3 and 3.3 significance levels with respect to the models of Radiodisk HESS0.32 and Radioprofile HESS0.32, respectively (see Table 1).
To be as completely as possible, we re-checked the extension of the counterpart of HESS J1616508 by setting Radiops for RCW 103. The uniform disk with a radius of 002–050 (in a 002 step) was tested again. We found that a 038 radius (thus the HESS0.38 model used above) resulted in a maximum value and the fit improvement of the model of Radiops HESS0.38 is at a 4.9 significance level with respect to that of Radiops HESS0.32 (Table 1). We thus considered this Radiops HESS0.38 model as the final one and used it in the following analysis.
Based on this model, we obtained 0.03, photon cm-2 s-1 for the counterpart of HESS J1616508 (TS=1233) and 0.07, photon cm-2 s-1 for RCW 103 (TS=165). Compared to the results previously reported in Xing et al. (2014), in which a point source was considered for HESS J1616508, the emission has a higher value (or softer). The -ray luminosity is 2.4 erg s-1 at a distance of 3.1 kpc.
2.4 Spectral Analysis
We extracted the -ray spectra of HESS J1616508 and RCW 103 by performing the maximum likelihood analysis of the LAT data in 10 evenly divided energy bands in logarithm from 0.1 to 500 GeV. In the extraction, the spectral normalizations of the sources within 5 degree of the center of HESS J1616508 were set as free parameters, while all the other parameters were fixed at the values obtained from the maximum likelihood analysis (the Radiops HESS0.38 model). Emissions of HESS J1616508 and RCW 103 were respectively described by a PL, with . For the obtained spectral data points, we kept those with TS9 and used the derived 95% flux upper limits instead when TS9 (Table 2).
We also evaluated the systematic uncertainties induced by the uncertainties of the Galactic diffuse emission. Following the commonly-used method (e.g., Abdo et al. 2009, 2010), we repeated the likelihood analysis in each energy band with the normalizations of the Galactic diffuse component artificially fixed to the 6% deviation values from the best-fit values. This 6% deviation represents the local departure from the best-fit diffuse model obtained from analyzing the source-free regions in the Galactic plane.
The obtained spectra with both statistical and systematic uncertainties are shown in Figure 3, while the flux and TS values of the spectral data points are provided in Table 2. The PL spectral models obtained for the two sources from the likelihood analysis match the spectral data points.
| Band (GeV) | TS | TS | ||
| 0.15 (0.1–0.2) | 1.43 | 0 | 1.73 | 0 |
| 0.36 (0.2–0.5) | 2.11 | 0 | 0.770.26 | 44 |
| 0.84 (0.5–1.3) | 1.170.41 | 119 | 0.720.63 | 67 |
| 1.97 (1.3–3.0) | 1.670.34 | 294 | 0.330.12 | 26 |
| 4.62 (3.0–7.1) | 1.430.26 | 198 | 0.370.07 | 47 |
| 10.83 (7.1–16.6) | 1.660.21 | 210 | 0.340.07 | 41 |
| 25.37 (16.6–38.8) | 1.630.21 | 146 | 0.18 | 1 |
| 59.46 (38.8–91.0) | 1.660.25 | 89 | 0.15 | 0 |
| 139.36 (91.0–213.3) | 1.560.33 | 44 | 0.28 | 0 |
| 326.60 (213.3–500.0) | 1.020.47 | 10 | 1.31 | 6 |
Note: is the energy flux () in 10-11 erg cm-2 s-1. Fluxes without uncertainties are the 95 upper limits.
2.5 Variability Check
We checked the -ray emission of RCW 103 for any long-term variability. Following the procedure introduced in Nolan et al. (2012), the variability index TSvar of a 60-day binned light curve (94 time bins) in 0.3–500 GeV was calculated, and we obtained TS. Given that a variable source would be identified at a 99% confidence level when TS127.6 (for 93 degrees of freedom), the emission did not show significant long-term variability.
3 Discussion
Using the latest Fermi-LAT source catalog, we have revisited the detection of -ray emission of RCW 103 by analyzing the data collected for years. The SNR’s emission is resolved at high-energy bands (Figures 1 & 2), even though the nearby source 4FGL J1616.25054e, the counterpart to the VHE source HESS J1616508, has been determined to be extended. Different from the previous result reported in Xing et al. (2014) that the emission was extended, we have found that it is consistent with being a PS. It has a PL spectrum with photon index , and this PL is softer than the previously reported ( in Xing et al. 2014). We note that this spectrum is very similar to that of the SNR Kes 17, which had (Supán et al., 2023).
The determined error-circle region of the PS emission coincides with the southern limb of RCW 103 (Figure 1), the brightest part of the SNR. No non-thermal X-ray emission has been detected from any part of RCW 103 (e.g., Braun et al. 2019), which otherwise may provide information for the origin of the -ray emission. On the other hand, multi-wavelength observations have shown different pieces of evidence for the interaction of the remnant with a molecular cloud (MC) in the vicinity of the southern limb (see, e.g., Oliva et al. 1999; Paron et al. 2006; Reach et al. 2006; Pinheiro Gonçalves et al. 2011; see also Braun et al. 2019 and references therein). Considering these, the positional coincidence thus strongly suggest a hadronic origin for the -ray emission.
We studied the broadband spectral energy distribution (SED) of the southern limb of RCW 103. The radio flux measurements of the whole SNR were provided in Beard (1966), Goss & Shaver (1970), Shaver & Goss (1970), Caswell et al. (1980), and Dickel et al. (1996), and we divided them by 2 to approximate the radio fluxes of the southern part. We analyzed archival X-ray data from one Chandra observation of RCW 103 (Obsid:18459), and obtained the X-ray flux in 0.5–7 keV from a region just outside the southern limb as the upper limit on non-thermal X-ray emission, where the value (unabsorbed) was erg s-1 cm-2. The radio fluxes and the X-ray flux upper limit, as well as the -ray fluxes, are shown in Figure 4.
To approximately fit the SED, we assumed that the particles accelerated in the SNR have a power-law form in energy with a high-energy cutoff:
where , the electrons or protons, is the normalization, and and are the power-law index and the cutoff energy, respectively. The total energy () is obtained from the integral of the equation. We used the Python package naima (Zabalza, 2015) to find reasonable fits to the SED.
By fitting the radio data with synchrotron radiation of the electrons, we obtained , magnetic field strength G, and erg, where was fixed at 1 TeV. The same population of the electrons also upper-scatter the low-energy photons in the field and produce -ray emission (i.e., the ICS process). Assuming the cosmic microwave background and the interstellar radiation field with a temperature of 30 K and an energy density of 0.5 eV cm-3 (Porter et al., 2006) at the source position, the produced -ray spectrum from ICS is shown in Figure 4, which is low compared to the observed -ray fluxes. It should noted that in this fitting practice, the allowed parameter ranges are much larger than those reported above, mainly because of the lack of a tight constraint in X-rays. Different combinations of the parameters can provide adequate fits to the radio data. However, we also noted that none of the combinations can simultaneously fit the -ray data. Because of the following hadronic model results (which have ), we chose to report the above parameters because of a similar electron PL index.
In order to fit the -ray data points with a hadronic model, in which -ray emission primarily arises from proton-proton collisions, we obtained and erg, where TeV was fixed and the target density was assumed to be (also adopted in Xing et al. 2014). As shown in Figure 4, the model spectrum can adequately describe the -ray data points. In addition, the spectrum drops to low flux values above TeV, which may explain the non-detection of this SNR in the TeV band (e.g., H. E. S. S. Collaboration et al. 2018). The hadronic-model fitting results, in combination with the positional coincidence of the -ray emission with the southern limb, thus point to the remnant-MC interaction as the likely origin for the high-energy emission from RCW 103.

### v2

@@
 Different templates for 4FGL J1616.25054e and RCW 103 are tested, and we find that a point source with a power-law (PL) spectrum at the southern limb of the SNR best describes the detected -ray emission.
-The photon index of the PL emission is , softer than the previously reported when the counterpart to HESS J1616518 was considered to be a point source (which likely caused mis-identification of extended emission at RCW 103).
+The photon index of the PL emission is (or in a Log-Parabola model), softer than the previously reported when the counterpart to HESS J1616518 was considered to be a point source (which likely caused mis-identification of extended emission at RCW 103).
 In order to produce the -ray emission in a hadronic scenario, we estimate that protons with an index2.4 PL energy distribution are needed.
@@
 0.3–500 GeV and 10–500 GeV as the representative ones, where the latter
-helps reveal the emission in a high, spatial resolution222http://www.slac.stanford.edu/exp/glast/groups/canda
-/lat_Perfomance.htm
+helps reveal the emission in a high, spatial resolution222https://www.slac.stanford.edu/exp/glast/groups/canda
+/lat_Performance.htm
 (smaller than the separation between HESS J1616508 and RCW 103).
@@
 | HESS0.40 | 43.2 |
-| CCO HESS0.32 | 57.1 |
+| PS HESS0.32 | 57.1 |
 | Radiodisk HESS0.32 | 63.5 |
@@
 of the central compact object (CCO) of RCW 103. The templates included
-a point source (PS; CCO model), a uniform disk with a radius of
+a point source (PS; PS model), a uniform disk with a radius of
 5 (given RCW 103’s size of 10 in diameter detected at
@@
 HESS0.32 template was still used. Performing the maximum likelihood
-analysis to the data with each of the models, CCO HESS0.32,
+analysis to the data with each of the models, PS HESS0.32,
 Radiodisk HESS0.32, and
@@
 Comparing the values, we found that the residual emission is more likely a point source located at the bright radio shell, as the fit improvement of the model Radiops HESS0.32 is at 4.3 and 3.3 significance levels with respect to the models of Radiodisk HESS0.32 and Radioprofile HESS0.32, respectively (see Table 1).
-To be as completely as possible, we re-checked the extension of the counterpart of HESS J1616508 by setting Radiops for RCW 103.
+To be as complete as possible, we re-checked the extension of the counterpart of HESS J1616508 by setting Radiops for RCW 103.
 The uniform disk with a radius of 002–050 (in a 002 step) was tested again.
@@
 The -ray luminosity is 2.4 erg s-1 at a distance of 3.1 kpc.
+We also tested a Log-Parabola model, , for the emission of RCW 103.
+The parameters and were set free, and , a scale parameter, was always fixed.
+Different values were tested.
+If we fixed it at 5 GeV, the resulting likelihood value indicated the fitting was better than that with the PL at a 2 significance level, where , , and photon cm-2 s-1 (TS ).
+The results are similar to those with the PL (note that the curvature parameter is very small), but the uncertainties are larger.
 2.4 Spectral Analysis
@@
 The PL spectral models obtained for the two sources from the likelihood analysis match the spectral data points.
+However for RCW 103, the two upper limits in 17–91 GeV (Table 2) are at the model-fit line, showing a slight inconsistency.
+Thus, its emission may be better described with the Log-Parabola model we obtained in Section 2.3, which is also shown in Figure 3.
 | Band (GeV) | TS | TS | ||
@@
 The SNR’s emission is resolved at high-energy bands (Figures 1 & 2), even though the nearby source 4FGL J1616.25054e, the counterpart to the VHE source HESS J1616508, has been determined to be extended.
-Different from the previous result reported in Xing et al.
+Differing from the previous result reported in Xing et al.
 (2014) that the emission was extended, we have found that it is consistent with being a PS.
-It has a PL spectrum with photon index , and this PL is softer than the previously reported ( in Xing et al.
+It has a PL spectrum with photon index or a Log-Parabola one with slope , and the spectrum is softer than the previously reported ( in Xing et al.
 2014).